OpenStudy (anonymous):

(w-6/w-8)-(w+1/w+8)+(w-40/w^2-64)=?????????

OpenStudy (mathteacher1729):

Are you being asked to simplify the following? \[\frac{w-6}{w-8}+\frac{w+1}{w+8}+\frac{w-40}{w^2-64}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the 1 st is minus

OpenStudy (anonymous):

(w-6/w-8)-(w+1/w+8)+(w-40/w^2-64)

OpenStudy (mathteacher1729):

\[\frac{w-6}{w-8}-\frac{w+1}{w+8}+\frac{w-40}{w^2-64}\]

OpenStudy (mathteacher1729):

Since you're being asked to simplify, you want all three fractions to have a common denominator. Luckily \[w^2-64 =(w+8)(w-8)\] so that means we can leave that term alone and multiply the others like this: \[\frac{w-6}{w-8}\cdot\frac{w+8}{w+8}-\frac{w+1}{w+8}\cdot\frac{w-8}{w-8}+\frac{w-40}{w^2-64}\] That LOOKS like a gigantic mess... But it gets better. We now have a common denominator... \[\frac{(w-6)(w-8) - (w+1)(w-8) + (w-40)}{w^2-64}\] And THAT just requires us to distribute and collect terms. :) Hope this helps!